OpenStudy (benlindquist):

The surface area of a rectangular prism is 108 centimeters squared. The length is 10 centimeters, and the width is 1 foot. What is the height of the prism?

OpenStudy (welshfella):

you have got mixed units there - centimeters and feet are you sure that is correct?

OpenStudy (benlindquist):

Let me check

OpenStudy (benlindquist):

16 centimeters 4.4 centimeters 4 centimeters 5.8 centimeters

OpenStudy (welshfella):

the width cant be 1 foot because 1 foot = about 30 cms so that would make the area of the base = about 300 cm2 and total surface area = 10 8

OpenStudy (welshfella):

i think thw width might be 1 cm

OpenStudy (benlindquist):

any other possible answers?

OpenStudy (welshfella):

lets assume width = 1 cm we have the diagram 

OpenStudy (benlindquist):

4.4 is not the right answer btw

OpenStudy (welshfella):

the total surface area = 2 * area of the base + 2 * 10*x + 2 * 1* x = 2*10 + 20x + 2x and the question says this equals 108 cm2 so we have the equation 20 + 20x + 2x = 108 22x + 20 = 108 solve this for x

OpenStudy (welshfella):

( x is the height of the prism)

OpenStudy (welshfella):

do you follow the above reasoning?

OpenStudy (benlindquist):

would you subtract 20 from 20 and 108 - 20 = ? and then divide 22x into what answer you got

OpenStudy (welshfella):

subtract 20 from both sides then divide by 22

OpenStudy (benlindquist):

22x = 86

OpenStudy (welshfella):

86???

OpenStudy (welshfella):

108 - 20 = ?

OpenStudy (benlindquist):

88

OpenStudy (welshfella):

right so 22x = 88 all you need to do now is divide 88 by 22

OpenStudy (anonymous):

hi hi hi hi hi hi hi hi

OpenStudy (benlindquist):

x = 25

OpenStudy (benlindquist):

the answer was 7??

OpenStudy (welshfella):

??? 88 / 22 = 4

OpenStudy (welshfella):

sounds like you are doing an online test..

OpenStudy (benlindquist):

oops, divided wrong.

OpenStudy (benlindquist):

its a lesson

OpenStudy (welshfella):

if 1 foot is really 1 cm the answer is 4 cm.

OpenStudy (benlindquist):

i divided wrong

OpenStudy (welshfella):

gotta go right now

OpenStudy (benlindquist):

ok